Could someone please help me to graph the line containing the given pair of points and find the slope?
(6, 3) (-5, -2)

Answer by mananth(16946)

x1 y1 x2 y2
6 3 -5 -2

slope m = (y2-y1)/(x2-x1) 0.45
m= 0.45

Plug value of the slope and point (6,3)in
Y =mx+b
3.00=2.73+b
b=3-2.73
b=0.27
So the equation will be
Y =0.45x+0.27